# Distilling Morphology-Conditioned Hypernetworks for Efficient Universal Morphology Control

### v1

@@ 7 Conclusion
In this paper, we propose HyperDistill to achieve both good performance and high efficiency at inference time for universal morphology control. While training HyperDistill via RL is hard, we take a policy distillation approach, and systematically investigate how some key algorithmic choices influence task generalization of the student policy. HyperDistill matches the performance of the universal TF teacher on both training and test robots, while significantly reducing memory and computational cost, which supports our hypothesis that decoupling inter-task and intra-task knowledge can improve inference efficiency. Given the wide application of TF in foundation models and their extremely high inference cost (Achiam et al., 2023; Brohan et al., 2022, 2023; Padalkar et al., 2023), the findings in our paper could serve as a general principle to reduce the inference cost of TF in other domains for potential future work.
Impact Statements
@@ B.3 Architecture Details of Different Methods
Table 2 shows the size of different architectures in each environment. For the TF teacher and TF (oracle), we use the same architecture hyperparameters as in Xiong et al. (2023). For the base MLP in HyperDistill, we fix the hidden layer size to 256, and do a grid search over hidden layer number to find the smallest base MLP that can achieve on-par performance with the teacher policy in each environment. Then we set MLP’s architecture identical to the base MLP in HyperDistill. For TF (compressed), as there is a two-layer MLP after each attention module in the standard TF implementation (Vaswani et al., 2017), we first reduce the dimension of this MLP’s hidden layer to 128, and reduce the number of attention heads to 1, then do a grid search over the number of attention layers to get a TF with a similar number of parameters as HyperDistill’s base MLP. The token embedding dimension is 128 for all different TF architectures.
Environment Base MLP in HyperDistill TF teacher TF (compressed)
& MLP & TF (oracle)
Layer num. Hidden size Layer num. Attention head Linear hidden size Layer num. Attention head Linear hidden size
FT 2 256 5 2 1024 1 1 128
VT 3 256 5 2 1024 2 1 128
Obstacle 3 256 5 2 1024 2 1 128
@@ Appendix C Further Experimental Results and Analysis
C.1 Main Results on the Training Robots
Figure 6 shows how different methods perform on the training robots during the distillation process. See Section 4.2 for analysis on the main results.
C.2 Further Analysis on Inference Efficiency
@@
Although MLP and the base network of HyperDistill have the same number of hidden layers and hidden units, the model size and FLOPs of MLP are still a bit larger than those of HyperDistill, as it needs to condition on the morphology context by concatenating limb-wise context features to the policy input, which introduces some additional cost.
C.3 Ablation Study on Context Representation Learning

### v2

@@ 7 Conclusion
In this paper, we propose HyperDistill to achieve both good performance and high efficiency at inference time for universal morphology control. While training HyperDistill via RL is hard, we take a policy distillation approach, and systematically investigate how some key algorithmic choices influence task generalization of the student policy. HyperDistill matches the performance of the universal TF teacher on both training and test robots, while significantly reducing memory and computational cost, which supports our hypothesis that decoupling inter-task and intra-task knowledge can improve inference efficiency.
Given the wide application of TF in foundation models and their extremely high inference cost (Achiam et al., 2023; Brohan et al., 2022, 2023; Padalkar et al., 2023), the findings in our paper could serve as a general principle to reduce the inference cost of TF in other domains as well. However, a potential limitation when extending our method to other domains is the additional cost introduced by the HN. This is not an issue for universal morphology control, as the morphology context remains unchanged on each robot. So we only need to call the HN once before deployment while the whole HN can be discarded afterwards. However, if task context changes over time, such as controlling a robot to solve different tasks by following language instructions, the HN can not be discarded and needs to be called whenever a new instruction is given. This requires additional space to save the HN parameters, and the inference efficiency gain may decrease as the HN will be called more frequently. How to tackle this challenge is an interesting direction for future work.
Impact Statement
@@
Acknowledgements
We would like to thank Matthew Jackson and Alex Goldie for their helpful discussion on the work. We would also like to thank the conference reviewers for their constructive feedback on the paper. Zheng Xiong is supported by UK EPSRC CDT in Autonomous Intelligent Machines and Systems (grant number EP/S024050/1) and AWS. Risto Vuorio is supported by EPSRC Doctoral Training Partnership Scholarship and Department of Computer Science Scholarship. Jacob Beck is supported by the Oxford-Google DeepMind Doctoral Scholarship. The experiments were made possible by a generous equipment grant from NVIDIA.
@@ B.3 Architecture Details of Different Methods
Table 2 shows the size of different architectures in each environment. For the ModuMorph (teacher) and ModuMorph (oracle), we use the same architecture hyperparameters as in Xiong et al. (2023). For the base MLP in HyperDistill, we fix the hidden layer size to 256, and do a grid search over hidden layer number to find the smallest base MLP that can achieve on-par performance with the teacher policy in each environment. Then we set multi-robot MLP’s architecture identical to the base MLP in HyperDistill. For ModuMorph (compressed) and TF (compressed), we tune the number of attention layers, the number of attention heads, and the dimension of the linear layer’s hidden dimension inside the attention module, so that the compressed model has a similar number of parameters as HyperDistill’s base MLP. The token embedding dimension is 128 for all different TF architectures.
Environment Base MLP in HyperDistill ModuMorph (teacher) ModuMorph (compressed) TF (compressed)
& Multi-robot MLP & ModuMorph (oracle)
Layer num. Hidden size Layer num. Head Hidden dim Layer num. Head Hidden dim Layer num. Head Hidden dim
FT 2 256 5 2 1024 1 1 128 1 1 256
VT 3 256 5 2 1024 1 1 128 2 1 128
Obstacle 3 256 5 2 1024 1 1 128 2 1 128
@@ Appendix C Further Experimental Results and Analysis
C.1 Further Analysis on Inference Efficiency
@@
Although multi-robot MLP and the base network of HyperDistill have the same number of hidden layers and hidden units, the model size and FLOPs of multi-robot MLP are still a bit larger than those of HyperDistill, as it needs to condition on the morphology context by concatenating limb-wise context features to the policy input, which introduces some additional cost.
In HyperDistill, generating the base MLP with HN takes 43M FLOPs in the FT environment, and 65M FLOPs in the VT and Obstacle environment. The FLOPs of generating the base MLP with HN is just slightly larger than the FLOPs of a single inference step with a universal TF controller.
C.2 Ablation Study on Context Representation Learning